# The HITRAN2024 methane update

### v2

@@ 1 Introduction
The spectroscopy of methane has a fundamental role in the observations and radiative transfer modeling of gaseous media, especially atmospheres of Earth [chanmiller_methane_2024], solar system planets and moons [10.1006/icar.1994.1139, 10.1016/j.icarus.2005.02.004], exoplanets (e.g. HD 102195b [10.1051/0004-6361/201834615]), brown dwarfs (e.g. HD 102195b [10.1093/MNRAS/STAC1412]), and combustion environments [10.1364/AO.35.004026]. The recent development of remote observation tools such as those on board Sentinel-5 [lorente_methane_2021], GOSAT, GOSAT-2 [imasu_greenhouse_2023], and MethaneSAT [chanmiller_methane_2024] has pushed the demand for higher quality parameters at many temperatures and pressures, in the presence of various other species, provided by the HITRAN database [gordon_hitran2020_2022]. Methane is frequently measured in planetary and stellar atmospheres, where its study plays an important role in understanding the formation and composition of these astronomical objects. Important methane-specific data such as its presence, concentration, or partial pressure would not be retrievable without models that incorporate accurate spectroscopic parameters.
@@
The octad (3500–5000 cm-1) and tetradecad regions (5000–6250 cm-1) in particular are crucial for satellite missions [lorente_methane_2021, malina_consistency_2022, chesnokova_estimation_2020], not only for its principal isotopologue but also for 13CH4 and CH3D, which are used to determine 13C/12C and D/H ratios. The octad has been improved in previous editions of HITRAN in the window observed by TROPOMI (4190–4340 cm-1). However, at higher wavenumbers of the octad, spectral parameters were of lower accuracy, incomplete, or both. Current and planned CH4 satellite sensors strive for  0.5% precision in the column (Ref. [jacob_quantifying_2022]; Table 1) for area flux mapping purposes, which requires similar accuracies for the intensities and widths and is not achieved in many regions. Even in the TROPOMI window, the HITRAN uncertainty code only reaches 4 (i.e., $0.0001\leq\Delta\nu<0.001$ cm-1 for line positions and $10\%\leq\frac{\Delta S_{ij}}{S_{ij}}<20\%$ for intensities).
@@
Above, and including the tetradecad region, other problems arise from the lack of assignment and/or a value for the lower state energy ($E^{\prime\prime}$). Given the temperature range, these issues are limited in the Earth’s atmosphere, but make accurate spectral modeling impossible at more extreme temperatures (for example, low temperatures on Titan or high temperatures on hot-Jupiters and brown dwarfs). With that one should note, that while for higher temperatures one can use the HITEMP version of the database [10.3847/1538-4365/ab7a1a], it also has limitations due to difficulty calculating accurate line positions at higher wavenumbers. Furthermore, the origin of some of the parameters in HITRAN2020 is not well documented, especially for pressure-induced data such as those being the results of private communications. This can cause additional problems, preventing comparisons, or extrapolations and adjustments if the old data was generated by a model.
@@
The focus of the HITRAN2024 update was on the parameters in the 160-characters HITRAN format, while a few speed-dependent and first-order line-mixing parameters were added as well. Significant progress has been made between 900 and 14000 cm-1 [kefala_empirical_2024, nikitin_improved_2024, rodina_improved_2021, nikitin_preliminary_2013, starikova_assignment_2024, malarich_dual_2021, nikitin_first_2024, campargue_high_2023] for the three most abundant isotopologues, 12CH4, 13CH4, and 12CH3D, allowing for an update to the transition wavenumber, intensity, and assignments from the HITRAN2020 database using various recorded spectra for validation. These line lists also cover missing portions of the infrared spectrum that were previously challenging to obtain, as well as some in-between lines that were missing. Similar efforts have been put into obtaining good quality pressure-induced line-shape parameters [antony_n2_2008, bertin_co2_2024, clark_difference_2004, devi_self-_2015, devi_spectral_2016, devi_multispectrum_2018, es-sebbar_intensities_2014, es-sebbar_linestrengths_2021, farji_airinduced_2021, gabard_line_2010, gharib-nezhad_h2induced_2019, ghysels_spectroscopy_2011, ghysels_temperature_2014, grigoriev_estimation_2001, lyulin_measurements_2014, ma_temperature_2016, manne_determination_2017, mondelain_line_2005, mondelain_measurement_2007, pine_speeddependent_2019, pine_speeddependent_2000, pine_multispectrum_2003, richard_self_2023, dudaryonok_semiempirical_2018], with the differences being that we considered older work, with the intention of replacing the parameters for more lines than those in Section 2 to improve traceability. To this end, a combination of good quality measurements, with an empirical Padé-approximants model based on these measurements and semi-empirical models to fill the gaps was used. These results provide the necessary foundation for the inclusion of first-order and full line-mixing with speed-dependent profiles.
@@ 2 Additions and updates of non-pressure-induced data
The newest versions of MARVELized-ExoMol [yurchenko_exomol_2024], TheoReTS [10.3847/1538-4357/aa8909], HITEMP [10.3847/1538-4365/ab7a1a] and MeCaSDa [MeCaSDa_2013], which is partly accessible via VAMDC portal [rixon_VAMDC_2011, albert_VAMDC_2020], line lists as well as lists from Nikitin et al. [nikitin_improved_2024, nikitin_preliminary_2013, nikitin_first_2024], Rodina et al. [rodina_improved_2021], Sung et al. [sung_new_2024], Starikova et al. [starikova_assignment_2024], Malarich et al. [malarich_dual_2021], and Campargue et al. [campargue_high_2023] were available for the HITRAN2024 update and are summarized for each region covered. These works were then validated against several high-quality, high-resolution FTS [2015ApJ...813...12H, bertin_co2_2024, birk_esa_2017, birk_measurement_2025] and CRDS spectra [birk_esa_2017] obtained under well-controlled laboratory conditions, when available for the updated region. The best parameters depending on the spectral region were selected and added to the HITRAN2024 database.
@@ 2.1.1 ExoMol
The ExoMol project [tennyson_2024_2024] has produced a number of methane line lists of increasing completeness and accuracy designed for use at high temperatures [yurchenko_exomol_2014, yurchenko_hybrid_2017, yurchenko_exomol_2024]. The intensities in this list are calculated variationally. The line positions in the most recent of these [yurchenko_exomol_2024], known as the MM line list, was based on a comprehensive MARVEL (measured active vibration-rotation energy levels) study of 12CH4 by Kefala et al [kefala_empirical_2024]. This MARVEL study extracted 82173 rovibrational and rotational transitions from 96 literature sources containing high-resolution studies of 12CH4 spectra; the study identified a further 42 sources containing unassigned or partially assigned methane spectra which could not be included in the MARVEL study. An analysis of these sources is currently in progress and has so far led to the assignment of almost 10000 further lines, which extend the coverage of the previous study in terms of coverage of both rotational and vibrational states. This work can enable updates to future releases of both HITRAN and HITEMP. A HITRAN-format 296 K version of the MM line list generated using PyExoCross [Zhang_pyExoCross_2024] was used in this study, predominantly in the pentad region.
@@ 2.1.2 TheoReTS
The TheoReTS database [rey_theorets_2016] provides line lists and quasi-continuum theoretical calculations for low and high temperatures from ab initio potential energy surfaces (PES) and dipole moment surfaces (DMS) together with spectra simulation tools. It was recently updated in [10.3847/1538-4357/aa8909] [rey_new_2018]. Variational line position calculations by Rey et al. have been refined using available empirical energy levels deduced from effective spectroscopic models. The TheoReTS predictions have been experimentally validated via high-temperature laser experiments by Ghysels et al. [ghysels_temp_2018] and by Wong et al. [Wong_atlas_2019] for their methane cross sections, as well as in dual frequency comb experiments and double resonance experiments by Malarich et al. [malarich_dual_2021] and Foltynowicz et al. [foltynowicz_2021] respectively. They were also used by Rey et al. [rey_new_2018] to model methane absorption at Titan, as observed by VIMS and by DISR instruments onboard the Cassini and Huygens space missions, and for the interpretation of non-LTE experiments in a hypersonic jet flow by Dudas et al. [dudas_2023]. In the context of the present HITRAN2024 update, TheoReTS predictions were used to create the HITEMP methane line list [10.3847/1538-4365/ab7a1a], while the extended assignments of 12CH4, 13CH4 and 12CH3D experimental spectra were carried out with improved line lists as described in the following sections. In recent multi-laboratory line intensity measurements, Reed et al. [reed_multilaboratory_2025] reported permille-level uncertainties of the R(0) to R(10) manifolds in the CH4 $2\nu_{3}$ band. These results had an average agreement of about 0.05 % with TheoReTS calculations which used the ab initio DMS of ref. [NIKITIN_dipole_intens_2017].
@@ 2.1.3 HITEMP
The HITEMP spectroscopic database [10.1016/j.jqsrt.2010.05.001] contains line lists for eight molecules, with the methane line list fully described in Hargreaves et al. [10.3847/1538-4365/ab7a1a]. The HITEMP list was constructed by combining the empirically corrected ab initio line list for 12CH4 from TheoReTS [10.3847/1538-4357/aa8909] with the HITRAN2016 line list [gordon_hitran2016_2017], which was the contemporary HITRAN list at the time. HITEMP CH4 spans 0-13,400 cm-1 and is suitable for simulating the high-temperature spectrum of methane up to 2000 K. In addition, the HITEMP list includes “effective” lines that are able to reproduce the continuum-like features provided by temperature-dependent lists of weak lines (i.e., so-called super-lines) in TheoReTS. Readers are referred to Hargreaves et al. [10.3847/1538-4365/ab7a1a] for a full description of the “effective” line methodology for methane.
@@ 2.1.4 MeCaSDa
CaSDa (Calculated Spectroscopic Databases) [RICHARD2024109127] is a project that presently gathers ten databases of line lists for ten highly symmetrical molecules including methane. These lists are built upon the analysis of high-resolution spectra (microwave, infrared, and Raman) using effective Hamiltonian and dipole moment operators written with a tensorial formalism [BOUDON2004620]. The databases also contain the full description of eigenvectors. Within this project, the MeCaSDa database was built for 12CH4 and 13CH4. The last update of MeCaSDa was performed in 2020 [RICHARD2020107096], following the results of a global fit described in Ref. [10.1063/1.5023331].
@@ 2.2.1 IAO FTIR measurements
Four spectra of methane samples at natural isotopic abundance were recorded between 900–1050 cm-1 at the Institute of Atmospheric Optics (IAO) in Tomsk [nikitin_improved_2024]. Nonlinear least-squares curve fitting procedures were used to retrieve line positions and intensities, which were subsequently analyzed using effective Hamiltonian and effective dipole moment models. The enhanced absorption at long paths allowed the measurement of new transitions: up to $J=28$ for cold bands and up to $J=20$ for hot bands. The new experimental list contains line positions and intensities for 2570 lines, with quantum assignments made for 1244 of these lines for the main isotopologue 12CH4. Comparisons of theoretical absorption simulations with experimental spectra revealed a considerable improvement compared to the HITRAN2020 database. All of the 1244 assigned line positions were fit with an average RMS standard deviation equal to 0.00065 cm-1.
@@ 2.2.2 HITRAN update in the 900–1050 cm-1 region
The new line list from Nikitin et al. [nikitin_improved_2024] (12CH4, 13CH4, 12CH3D & 13CH3D) completes the lower energy part of the dyad region, adding many lines including those for the lower-abundance 13CH3D isotopologue. Their data were merged directly to the region, replacing the transition wavenumbers for any difference ($\Delta\tilde{\nu}=0$ cm-1) between HITRAN2020 and the new parameters, the lower energy levels for $\Delta E^{\prime\prime}=0.1$ cm-1, and the intensities for $\Delta S_{ij}/S_{ij}=0.1$. About half of the lines in the list mainly served to update the positions and intensities with a few energy levels and assignments, while the other half was entirely new.
@@ 2.3.1 HITRAN update in the 2500–3900 cm-1 region
The pentad region is relatively well defined in HITRAN2020, nevertheless, we have incorporated a few noticeable improvements. Spectra from Hargreaves et al. [2015ApJ...813...12H] ($S1$) and Bertin & Vander Auwera [bertin_co2_2024] ($S2$) were used to validate the spectral parameters. Although $S1$ spans a much larger region (2500–4900 cm-1), $S2$ (2860–3160 cm-1) provides a higher signal-to-noise ratio (SNR) and higher resolution, making it a better choice for validating the line list in this window. Small missing parts appear as well in $S1$ from cutting water absorption; making these small windows unusable for validation.
Figure 1: Residuals obtained at 296 K between experiment from $S2$ Bertin & Vander Auwera [bertin_co2_2024] and Voigt calculations from the line lists used for the HITRAN2024 update
@@ 2.4.1 Kitt Peak & JPL FTS measurements of 13CH4
Multiple sets of 13CH4 line lists, either empirical or model predictions, were available from various sources, e.g., HITRAN 2020 [gordon_hitran2020_2022, brown_measurements_2016], JPL GFIT package (“atm.161”) [toon_spectrometric_2021], and Kitt Peak FTS (291 K) and JPl-FTS (80 K). More importantly, comprehensive updates to the previous work by Brown et al. [brown_measurements_2016] became available through the effective Hamiltonian modeling extended to the tetradecad region for hot band transitions. All of the line lists have been evaluated and cross-compared against high-resolution laboratory spectra obtained at 291 K and 80 K by generating synthetic spectra for the the same experimental conditions (e.g. sample pressure, temperature, gas cell pathlength, and spectral resolution) as the corresponding laboratory spectra, but using different sets of line lists available. These multiple sets of synthetic spectra were compared with lab spectra at each temperature to determine their quantum identifications and lower state energy values by adopting model calculation as presented in Sung et al. [sung_new_2024]
@@
Coy, Bourssard, Toon, Sung [sung_new_2024] 4000–4700 cm-1 $1\times 10^{-26}$ cm-1atm-1 7103 / 6585
Table 1 presents statistics on the improvements (both additions and updates inclusive) as reported by Sung et al. [sung_new_2024], showing an almost ten-fold increase in the total number of line list entries in the 4000–4700 cm-1 wavenumber interval for 13CH4, when compared to the HITRAN2020.

### v3

@@ 1 Introduction
The spectroscopy of methane has a fundamental role in the observations and radiative transfer modeling of gaseous media, especially atmospheres of Earth [14], solar system planets and moons [52, 50], exoplanets (e.g. HD 102195b [40]), brown dwarfs (e.g. HD 102195b [106]), and combustion environments [66]. The recent development of remote observation tools such as those on board Sentinel-5 [55], GOSAT, GOSAT-2 [49], and MethaneSAT [14] has pushed the demand for higher quality parameters at many temperatures and pressures, in the presence of various other species, provided by the HITRAN database [37]. Methane is frequently measured in planetary and stellar atmospheres, where its study plays an important role in understanding the formation and composition of these astronomical objects. Important methane-specific data such as its presence, concentration, or partial pressure would not be retrievable without models that incorporate accurate spectroscopic parameters.
@@
The octad (3500–5000 cm-1) and tetradecad regions (5000–6250 cm-1) in particular are crucial for satellite missions [55, 61, 15], not only for its principal isotopologue but also for 13CH4 and CH3D, which are used to determine 13C/12C and D/H ratios. The octad has been improved in previous editions of HITRAN in the window observed by TROPOMI (4190–4340 cm-1). However, at higher wavenumbers of the octad, spectral parameters were of lower accuracy, incomplete, or both. Current and planned CH4 satellite sensors strive for  0.5% precision in the column (Ref. [51]; Table 1) for area flux mapping purposes, which requires similar accuracies for the intensities and widths and is not achieved in many regions. Even in the TROPOMI window, the HITRAN uncertainty code only reaches 4 (i.e., $0.0001\leq\Delta\nu<0.001$ cm-1 for line positions and $10\%\leq\frac{\Delta S_{ij}}{S_{ij}}<20\%$ for intensities).
@@
Above, and including the tetradecad region, other problems arise from the lack of assignment and/or a value for the lower state energy ($E^{\prime\prime}$). Given the temperature range, these issues are limited in the Earth’s atmosphere, but make accurate spectral modeling impossible at more extreme temperatures (for example, low temperatures on Titan or high temperatures on hot-Jupiters and brown dwarfs). With that one should note, that while for higher temperatures one can use the HITEMP version of the database [42], it also has limitations due to difficulty calculating accurate line positions at higher wavenumbers. Furthermore, the origin of some of the parameters in HITRAN2020 is not well documented, especially for pressure-induced data such as those being the results of private communications. This can cause additional problems, preventing comparisons, or extrapolations and adjustments if the old data was generated by a model.
@@
The focus of the HITRAN2024 update was on the parameters in the 160-characters HITRAN format, while a few speed-dependent and first-order line-mixing parameters were added as well. Significant progress has been made between 900 and 14000 cm-1 [53, 75, 95, 76, 100, 60, 72, 12] for the three most abundant isotopologues, 12CH4, 13CH4, and 12CH3D, allowing for an update to the transition wavenumber, intensity, and assignments from the HITRAN2020 database using various recorded spectra for validation. These line lists also cover missing portions of the infrared spectrum that were previously challenging to obtain, as well as some in-between lines that were missing. Similar efforts have been put into obtaining good quality pressure-induced line-shape parameters [3, 6, 17, 21, 22, 20, 25, 26, 27, 30, 33, 34, 35, 39, 58, 59, 62, 63, 65, 81, 79, 80, 91, 23], with the differences being that we considered older work, with the intention of replacing the parameters for more lines than those in Section 2 to improve traceability. To this end, a combination of good quality measurements, with an empirical Padé-approximants model based on these measurements and semi-empirical models to fill the gaps was used. These results provide the necessary foundation for the inclusion of first-order and full line-mixing with speed-dependent profiles.
@@ 2 Additions and updates of non-pressure-induced data
The newest versions of MARVELized-ExoMol [124], TheoReTS [89], HITEMP [42] and MeCaSDa [4], which is partly accessible via VAMDC portal [94, 1], line lists as well as lists from Nikitin et al. [75, 76, 72], Rodina et al. [95], Sung et al. [102], Starikova et al. [100], Malarich et al. [60], and Campargue et al. [12] were available for the HITRAN2024 update and are summarized for each region covered. These works were then validated against several high-quality, high-resolution FTS [41, 6, 7, 8] and CRDS spectra [7] obtained under well-controlled laboratory conditions, when available for the updated region. The best parameters depending on the spectral region were selected and added to the HITRAN2024 database.
@@ 2.1.1 ExoMol
The ExoMol project [107] has produced a number of methane line lists of increasing completeness and accuracy designed for use at high temperatures [126, 125, 124]. The intensities in this list are calculated variationally. The line positions in the most recent of these [124], known as the MM line list, was based on a comprehensive MARVEL (measured active vibration-rotation energy levels) study of 12CH4 by Kefala et al [53]. This MARVEL study extracted 82173 rovibrational and rotational transitions from 96 literature sources containing high-resolution studies of 12CH4 spectra; the study identified a further 42 sources containing unassigned or partially assigned methane spectra which could not be included in the MARVEL study. An analysis of these sources is currently in progress and has so far led to the assignment of almost 10000 further lines, which extend the coverage of the previous study in terms of coverage of both rotational and vibrational states. This work can enable updates to future releases of both HITRAN and HITEMP. A HITRAN-format 296 K version of the MM line list generated using PyExoCross [127] was used in this study, predominantly in the pentad region.
@@ 2.1.2 TheoReTS
The TheoReTS database [83] provides line lists and quasi-continuum theoretical calculations for low and high temperatures from ab initio potential energy surfaces (PES) and dipole moment surfaces (DMS) together with spectra simulation tools. It was recently updated in [89] [84]. Variational line position calculations by Rey et al. have been refined using available empirical energy levels deduced from effective spectroscopic models. The TheoReTS predictions have been experimentally validated via high-temperature laser experiments by Ghysels et al. [36] and by Wong et al. [121] for their methane cross sections, as well as in dual frequency comb experiments and double resonance experiments by Malarich et al. [60] and Foltynowicz et al. [28] respectively. They were also used by Rey et al. [84] to model methane absorption at Titan, as observed by VIMS and by DISR instruments onboard the Cassini and Huygens space missions, and for the interpretation of non-LTE experiments in a hypersonic jet flow by Dudas et al. [24]. In the context of the present HITRAN2024 update, TheoReTS predictions were used to create the HITEMP methane line list [42], while the extended assignments of 12CH4, 13CH4 and 12CH3D experimental spectra were carried out with improved line lists as described in the following sections. In recent multi-laboratory line intensity measurements, Reed et al. [82] reported permille-level uncertainties of the R(0) to R(10) manifolds in the CH4 $2\nu_{3}$ band. These results had an average agreement of about 0.05 % with TheoReTS calculations which used the ab initio DMS of ref. [78].
@@ 2.1.3 HITEMP
The HITEMP spectroscopic database [97] contains line lists for eight molecules, with the methane line list fully described in Hargreaves et al. [42]. The HITEMP list was constructed by combining the empirically corrected ab initio line list for 12CH4 from TheoReTS [89] with the HITRAN2016 line list [38], which was the contemporary HITRAN list at the time. HITEMP CH4 spans 0-13,400 cm-1 and is suitable for simulating the high-temperature spectrum of methane up to 2000 K. In addition, the HITEMP list includes “effective” lines that are able to reproduce the continuum-like features provided by temperature-dependent lists of weak lines (i.e., so-called super-lines) in TheoReTS. Readers are referred to Hargreaves et al. [42] for a full description of the “effective” line methodology for methane.
@@ 2.1.4 MeCaSDa
CaSDa (Calculated Spectroscopic Databases) [93] is a project that presently gathers ten databases of line lists for ten highly symmetrical molecules including methane. These lists are built upon the analysis of high-resolution spectra (microwave, infrared, and Raman) using effective Hamiltonian and dipole moment operators written with a tensorial formalism [9]. The databases also contain the full description of eigenvectors. Within this project, the MeCaSDa database was built for 12CH4 and 13CH4. The last update of MeCaSDa was performed in 2020 [92], following the results of a global fit described in Ref. [2].
@@ 2.2.1 IAO FTIR measurements
Four spectra of methane samples at natural isotopic abundance were recorded between 900–1050 cm-1 at the Institute of Atmospheric Optics (IAO) in Tomsk [75]. Nonlinear least-squares curve fitting procedures were used to retrieve line positions and intensities, which were subsequently analyzed using effective Hamiltonian and effective dipole moment models. The enhanced absorption at long paths allowed the measurement of new transitions: up to $J=28$ for cold bands and up to $J=20$ for hot bands. The new experimental list contains line positions and intensities for 2570 lines, with quantum assignments made for 1244 of these lines for the main isotopologue 12CH4. Comparisons of theoretical absorption simulations with experimental spectra revealed a considerable improvement compared to the HITRAN2020 database. All of the 1244 assigned line positions were fit with an average RMS standard deviation equal to 0.00065 cm-1.
@@ 2.2.2 HITRAN update in the 900–1050 cm-1 region
The new line list from Nikitin et al. [75] (12CH4, 13CH4, 12CH3D & 13CH3D) completes the lower energy part of the dyad region, adding many lines including those for the lower-abundance 13CH3D isotopologue. Their data were merged directly to the region, replacing the transition wavenumbers for any difference ($\Delta\tilde{\nu}=0$ cm-1) between HITRAN2020 and the new parameters, the lower energy levels for $\Delta E^{\prime\prime}=0.1$ cm-1, and the intensities for $\Delta S_{ij}/S_{ij}=0.1$. About half of the lines in the list mainly served to update the positions and intensities with a few energy levels and assignments, while the other half was entirely new.
@@ 2.3.1 HITRAN update in the 2500–3900 cm-1 region
The pentad region is relatively well defined in HITRAN2020, nevertheless, we have incorporated a few noticeable improvements. Spectra from Hargreaves et al. [41] ($S1$) and Bertin & Vander Auwera [6] ($S2$) were used to validate the spectral parameters. Although $S1$ spans a much larger region (2500–4900 cm-1), $S2$ (2860–3160 cm-1) provides a higher signal-to-noise ratio (SNR) and higher resolution, making it a better choice for validating the line list in this window. Small missing parts appear as well in $S1$ from cutting water absorption; making these small windows unusable for validation.
Figure 1: Residuals obtained at 296 K between experiment from $S2$ Bertin & Vander Auwera [6] and Voigt calculations from the line lists used for the HITRAN2024 update
@@ 2.4.1 Kitt Peak & JPL FTS measurements of 13CH4
Multiple sets of 13CH4 line lists, either empirical or model predictions, were available from various sources, e.g., HITRAN 2020 [37, 10], JPL GFIT package (“atm.161”) [108], and Kitt Peak FTS (291 K) and JPl-FTS (80 K). More importantly, comprehensive updates to the previous work by Brown et al. [10] became available through the effective Hamiltonian modeling extended to the tetradecad region for hot band transitions. All of the line lists have been evaluated and cross-compared against high-resolution laboratory spectra obtained at 291 K and 80 K by generating synthetic spectra for the the same experimental conditions (e.g. sample pressure, temperature, gas cell pathlength, and spectral resolution) as the corresponding laboratory spectra, but using different sets of line lists available. These multiple sets of synthetic spectra were compared with lab spectra at each temperature to determine their quantum identifications and lower state energy values by adopting model calculation as presented in Sung et al. [102]
@@
Coy, Bourssard, Toon, Sung [102] 4000–4700 cm-1 $1\times 10^{-26}$ cm-1atm-1 7103 / 6585
Table 1 presents statistics on the improvements (both additions and updates inclusive) as reported by Sung et al. [102], showing an almost ten-fold increase in the total number of line list entries in the 4000–4700 cm-1 wavenumber interval for 13CH4, when compared to the HITRAN2020.